Oxford Materials IT Committee Meeting


There will be a meeting at 11am on Tuesday 31st January 2017 in Banbury Road Committee Room.


  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters outstanding from the previous meeting
  • 3 Matters for discussion (attached)
  • 4 Head of IT report (attached)
  • 5 Finances
  • 6 AOB
  • 7 Date and time of next meeting: Tuesday 9th May 2017 at 11am in BR Committee Room.


See minutes below.


Verbal reports on the status of various projects:

IT services projects: Chorus, OpenAccess, SecurityBaseline
Departmental projects: Server infrastructure, Photocopiers, iLab, iProcurement
Information from other Committees: MPLS ICT Panel, Divisional Software Licensing Group


The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Personnel: The IT review report was never supplied. The Grade 6 IT staff position remains on hold but was outside the department’s reported headcount in December 2015 so is unlikely to be released. PJW worked up draft plans for possible relocation of IT support area to HR, but notes the plan will separate office from workshop activities.

Admin Network Lots of staffing changes again. New finance staff Kiri Holmes, Raya Batishteva and new receptionist Tara Haydon all joined in the New Year. Jane Foxall started as GADB PA

Teaching Laboratory All computers were checked/patched before the start of term. The two week Modelling option course was attended by 22 students using personal laptops, but the six laptops purchased last July and the six teaching lab laptops were used for using COMSOL.

Digital Signage: Electronic display signage continues to be updated by Lorraine Laird.

Lecture theatres / meeting rooms New roller whiteboards were fitted by IT staff in HRLT and ETBCR. New roller whiteboard for BRLT has been ordered.

Printers and photocopiers: Annual papercut license renewal is still outstanding from July. PJW has taken management of photocopiers and toner ordering back from reception staff.

Digital Print-room No changes.

Library Both library workstations are in need of replacement with trickledown hardware.

MML There are still no funds to repair the MML aircon which is running at 50% capacity but stable since it is winter. Hahn has been mended and reinstalled. MMLfarm repaired hardware. Landau was reinstalled after a disk failure. Kittel is stable. Ironman still has broken nodes and is in need of reinstallation. Quotes were obtained to support a grant application to start replacing Ironman.

FIM Dan Haley has taken over from Tomas Martin as IT liaison. No significant changes.

EM Consolidation of EM activities to central site has finished. Problems with PDNGPU and PDNHPC resolved.

OMCS No significant changes.

MFFP Monthly patching and rebooting of 13 systems and storage arrays. MFFP website is desperate for upgrading due to Drupal6 end of life last April. UPS protection of storage arrays was reorganised. AJW purchased more storage. TJM is purchasing more storage. Diskspace cleared on several workstations.

PGB Salto door entry system still needs fixing to work with university cards rather than separate cards. Steady stream of new people and equipment. Disk fail on XRD control PC.

Website Daily/weekly editing of department website and news items. More than 30 still need migrating, mostly due to CmsMadeSimple needing upgrading 1.13 to 2.1 and configuring for https and where possible incorporating responsive design. MFFP website needs migrating to Drupal7. The university Mosaic CMS is still not open to new users. All 20 SSL certificates were renewed.

Publications Weekly editing of Departmental Newsletter. Only 4 sets of business cards. Signage for Hirsch Lecture. Posterboards for workshops (Oxford ECS, QIP). Printed handbooks for Finals. Dept handbook still pending.

Network Still no new network connections can be made within the Hume-Rothery building (or until a new route is created avoiding asbestos).

Security Oxcert released 7 security bulletins but none since December. Several minor virus infections dis-infected. The university continues to be targeted with phishing emails, and one academic was caught leading to SSO account compromise. One cease&desist copyright infringement notice was dealt with. The new windows update server now has 575 devices, whereas only 30 active devices remain on the old server.

WDE Still only six user’ devices encrypted – need to expand uptake targeting senior staff. Limited by time/manpower….

Disposal Currently collecting equipment again for another disposal van load.

Software COMSOL and Mathematica and Labview were renewed.

Disaster Recovery and Backup Aircon in the Hume-Rothery server room was replaced by estates department. PJW is still looking out for a second-hand replacement Symmetra UPS power module (<£1200) – currently this UPS is operating below capacity but is an acceptable risk. PJW still needs to finish rolling out purchased research group UPS’s into production (KP1/KP2/KP3/KAQOR).

Documentation updates Documentation on the private Mediawiki continues to expand, but time for documentation has been limited due to being short staffed.

New people & accounts Nov 4; Dec 8; Jan 9;

New Device registrations Nov 24; Dec 7; Jan 14;


A meeting of the Department of Materials IT Committee was held 11am on Tuesday 31st January 2017 in Banbury Road Lecture Theatre (moved from the BR Committee Room).

The following people attended: CS, PJW, RS, CA, SLP, JM, LB, Lewys Jones, Sabin Sulzer.

1. Apologies for Absence and Membership of the Committee

Flynn Castles, Andy Naylor and Richard Turner sent their apologies.

Paul Warren welcomed both Jan Mol (JM) and Lapo Bogani (LB) to the committee, both appointed by PSG replacing Paul Bagot. Also Sabin Sulzer was welcomed as JCCG chairman, replacing Jiale Wang.

2. Matters arising from the minutes

(Items marked with * were not for discussion at the meeting)

a) *Draft Information Security Policy (Minute 2a)

PJW still needs to implement a departmental Information Security policy, supported by PSG

Action: PJW/PSG

b) Infrastructure outage risks (Minute 2c)

After the outage in May 2016, APD requested that IT staff put together a document detailing the main areas of risk associated with single points of failure within departmental IT systems, and presenting costed proposals on how to minimize future risks. This document could then feed into the forthcoming Review of IT by IT services.

The requested document was not produced and the IT review has happened (but not reported). IT staff have attended a VMware training course (PJW/RS) and replacement switching for the VM infrastructure is being implemented to address performance issues. It is still desirable to have a document ready for consideration for budget planning in March.


c) Print cost reporting  (Minute 2e)

A newsletter item was written promoting people to reduce paper usage and advertising the login to the web interface at https://print.materials.ox.ac.uk:9192/user to view usage history.

d) Display Signage (Minute 2h)

SLP requested that the departmental display signage system be extended into the Holder Building cafeteria in order to counter-balance the Engineering display signage. IT staff have considered this request and identified a possible location above the chiller cabinet, however this expenditure (~£500) was not budgeted for this academic year and funds are tight. Instead CS has arranged for Materials to have four slides on the Engineering display signage, as supplied by LL who updates the Materials display signage.

e) Begbroke Map  (Minute 2i)

PJW updated the Begbroke map as used in the visitor travel leaflet and in the handbooks.

f) Newsletter item about “At Risk” period  (Minute 3)

PJW supplied LL with a Newsletter item (30th October 2016) informing everyone about the IT Services “At Risk” period of 7-9am each Tuesday when online services may be unavailable due to maintenance tasks.

g) Wireless in Begbroke Annex  (Minute 4)

IT staff installed a wireless access point in the Begbroke Annex in liaison with RT.

h) Photocopier management  (Minute 6)

Due to several instances of photocopiers running out of toners, PJW has taken ownership of monitoring the photocopiers and ordering replacement toners. This task may return to being managed by the new receptionist Tara Haydon in due course.

3 Matters for Discussion

PJW updated the committee verbally upon the status of the following.

IT services projects:

  • Chorus – the new telephone system http://projects.it.ox.ac.uk/icp/ Also see official Chorus newsletters http://projects.it.ox.ac.uk/icp/newsletter PJW is completing the network readiness questionnaire for each of our 13 buildings and managing the pre-migration data collection and validation for each telephone line (handsets/phonenumbers/people/directoryentries/networkports/powerplugs). Most Materials buildings are scheduled for the week commencing 8th May, but Holder Building will be done in the week commencing 22nd May. Rex Richards schedule is still under review.
  • OpenAccess http://openaccess.ox.ac.uk/ PSG has appointed Chris Akinola to manage the checking that papers have been deposited correctly, including checking dates of acceptance. PSG and PJW updated academics about the overall compliance status at the recent Away Day meeting. The system still depends on all researchers uploading as soon as the paper is accepted, because the bibliographic databases may not list the publication until after the 3 month deadline.
  • Security Baseline Questionnaire https://www.infosec.ox.ac.uk/guidance-policy/assurance The HoD has received the Security Baseline Questionnaire. The aim of the questionnaire is to gather a few of current status rather than to fix everything immediately. PJW has completed the technical questions last week and the high level review is being completed this week prior to submission later by CS.

    Action: PJW/CS/PSG

  • Information Security Improvement Programme (ISIP) This department still needs to review and adopt an Information Security Policy.(Minute 2a). Completing the Security Baseline Questionnaire has highlighted several areas where the department does not yet fully comply with the University Information Security Policy. PJW noted that back in 2012 when he was appointed DISO he requested some support resource from DC to document and implement the Information Security Policy, but this was not forthcoming. Also an IT support position incorporating responsibilities for Information Security was interviewed in 2015 but unfilled and then fell foul of the head count cap. PJW to discuss again with CS and PSG.

    Action: PJW/CS/PSG

    PJW noted that since July 2016 the Online Awareness Module has become compulsory for all staff and should be embedded in the staff induction processes and should be repeated annually. PJW presented a summary of who has not yet completed the awareness module. ITC agreed to formally raise with DC the problem of non-compliance despite repeated reminders. Currently there is no negative effect on staff if they ignore the request to complete the online awareness module. After some discussion it was agreed to propose that access to printers/photocopiers should be revoked if staff do not comply by 5th March. (which is when the next compliance report will be available). ITC also agreed that the Awareness Module should be recommended for all postgraduate students.

    Action: PJW->DC

  • Replay (automatic lecture recording service) http://help.it.ox.ac.uk/replay/index The Replay service has ended pilot and entered production. At present it is still free. Several departments now routinely record all their lectures. This department could follow that route but would need video (to cover the chalk&talk) in addition to the baseline audio+powerpoint. Potential for Begroke remote attendance to lectures. PJW discussed this with AOT and will add suitable video cameras for HRLT and BRLT to the March budget proposal.

    Action: PJW

  • Nexus Upgrade Office365 IT services are moving forward with the plan to replace Nexus Email service with Office365 online service (and 50Gb quotas) due to start in September 2017. There is evidence that the preparatory works have been progressing, however September is approaching rapidly.
  • Electronic Lab Notebooks IT services are currently assessing whether a centrally supported ELN solution would be beneficial. All Oxford researchers, lab managers and research support staff are invited to complete the ELN survey. Newsletter item 27 January 2017.

Departmental projects:

  • Server infrastructure Server infrastructure – IT staff are currently testing new switching for the VM infrastructure to address on-going latency problems, prior to scheduling some downtime to replace the switching. IT staff are obtaining quotes for storage to reduce the risk of single points of failure to feed into the budget proposal in March.

    Action: RS/PJW

  • Photocopiers Next year’s budget should contain a line for a copier to 14 Parks Road.

    Action: PJW

  • iLab CS reported that Neil Young is due to roll out training to group leaders such that they are better equipped to monitor the usage/cost of their group’s activities via iLab. PJW will check with Neil about updating the existing iLab training web information.

    Action: PJW

  • R12 iProcurement – CS reported that Debbie Townsend is starting a campaign to widen the use of R12 for ordering, aiming to depreciate the use of paper requisitions.

Information from other committees:

  • MPLS ICT panel – the Hilary Term meeting yesterday received a status report from IT Services, mostly all covered by the online Hilary Term newsletter. Items of interest have all been covered above. http://www.it.ox.ac.uk/newsletters/issue10
  • Divisional Software Licensing Group http://help.it.ox.ac.uk/sls/background The registration software webpage https://register.it.ox.ac.uk/software is now being used to distribute addition site licensed software such as SPSS, ArcGIS and NVivo as well as Matlab, CiscoAnyConnectVPN, Oxygen and Sophos. DSLG have triggered a review of antivirus options since Sophos license is due for renewal in ~18months.

4. Head of IT Report
A written report was circulated with the agenda.

The compromise of an SSO account was discussed. The likely trajectory was the use of an unsecure network on a mobile device. It was suggested to strongly recommend the use of VPN encryption if using open wireless networks.

Action: PJW

LJ and SLP questioned IT staff about the checking of computers in the practical labs. IT staff have a monthly cycle of applying security updates and checking operation. Over the past year there have been a few instances of the software not performing as expected during practical classes after system updates and software upgrades. It was agreed that it must be the responsibility of Demonstrators to check that the practical equipment, including the computers, is operational prior to the practical.

IT staff noted that several practical rely on old unsupported software, some of which limits the computers to only running 32bit version of the operating system, whereas many companies are moving towards dropping support for 32bit versions on their software.

It was agreed that IT staff should discuss these issues with Paula Topping and request that Senior Demonstrators review the software requirements of their practicals with a view to moving to 64bit OS support.

Action: PJW/RS

5. Finances
No status report was presented.

 6. AOB

There was no other business.

7. Date of Next Meeting

The next meeting is Tuesday 9th May 2017 at 11am in BR Committee Room.

The meeting closed at 12.30pm.